# Orchestrated Robust Controller for the Precision Control of Heavy-duty Hydraulic Manipulators

### v2

@@ Abstract
Vast industrial investment along with increased academic research on hydraulic heavy-duty manipulators has unavoidably paved the way for their automatization, necessitating the design of robust and high-precision controllers. In this study, an orchestrated robust controller is designed to address the mentioned issue for a generic manipulator with an anthropomorphic arm and spherical wrist. To do so, the entire robotic system is decomposed into subsystems, and a robust controller is designed at each local subsystem by considering unknown model uncertainties, unknown disturbances, and compound input nonlinearities, thanks to virtual decomposition control (VDC). As such, radial basic function neural networks (RBFNNs) are incorporated into VDC to tackle unknown disturbances and uncertainties, resulting in novel decentralized RBFNNs. All robust local controllers designed at each local subsystem, then, are orchestrated to accomplish high-precision control. In the end, for the first time in the context of VDC, a semi-globally uniformly ultimate boundedness is achieved under the designed controller. The validity of the theoretical results is verified by performing extensive simulations and experiments on a 6-degrees-of-freedom industrial manipulator with a nominal lifting capacity of $600\,kg$ at $5$ meters reach. Comparing the simulation result to the state-of-the-art controller along with provided experimental results, demonstrates that proposed method established all promises and performed excellently.
@@
SGUUB semi-globall UUB
@@ I-A Heavy-duty Hydraulic Manipulators
Because of their higher robustness and remarkably larger payload-to-weight ratio (PWR), heavy-duty hydraulic manipulators (HHMs) have been widely utilized in various fields, including construction, forestry, and agriculture. For example, the commercial manipulator shown in Fig. 1 has a lifting capacity of 600 kg at 5 meters reach, with a total weight of 445 kg (PWR almost 1:1), whereas the ABB robot [1] has a lifting capacity of 550 kg at 4.2 meters with the robot weight of 4,600 kg (PWR of 1:8). Such structural rigidity of the electrical manipulators is important to achieve higher accuracy in the manufacturing industry, while in the field robotic applications, lightweight HHMs are particularly suited for installation on mobile machines, addressing the critical requirement for traversability. Consequently, it has become crucial to design a high-precision controller for HHMs that preserves their superior traversability feature while attaining the precision of electrical manipulators. The current vast amount of industrial investments with increased academic interests will revolutionize the HHM industry [2, 3, 4, 5].
Despite their advantages, HHMs suffer from fundamental challenges, such as model nonlinearities and complexities emerging from structural complicatedness and fluid dynamics[5]. In contrast to electric actuators, whose model complexity can be neglected, hydraulic actuators are governed by fluid dynamics and may be subjected to nonsmooth and discontinuous nonlinearities due to friction between cylinder and piston, compound input nonlinearities, and valve under/overlap. Opposite to the electrically actuated open-chain manipulators, HHMs mostly comprise parallel-serial structures [6] that convert linear piston motion to rotational motion. Meanwhile, in rotary hydraulic actuators, the helical gears are employed to convert the piston motion to rotation, the same mechanism in the wrist actuators of the manipulator in this study (Fig. 1), one of which can deliver $2,000\,N\cdot m$ of torque. The mentioned nonlinearities and complexities render the stability analysis and control design of HHMs significantly challenging, emphasizing the need for robust and high-performance control schemes, not only in theory but also in real-world applications.
@@ I-B1 Model-based controller
In [8], different NMBC methods, such as sliding mode control (SMC), adaptive inverse dynamics controller (AIDC), and model-reference adaptive controllers with velocity measurement (MRACV), have been designed and compared, unveiling the better performance of MRACV in comparison to AIDC and SMC. In [9], an observer-based adaptive robust controller (ARC) has been designed to control a hydraulic manipulator, in which the backstepping approach has been employed to derive the control law. However, the experimental results in [10] with desired compensation ARC (DCARC) have shown that DCARC yields lower tracking error than ARC. Further, in [11] an adaptive output force tracking controller has been designed to control a 6-degrees-of-freedom (DoF) hydraulic manipulator by employing the virtual decomposition control (VDC) scheme, which was first introduced in [12]. A VDC control, also, in [13] has been designed for the precision control of a hydraulic crane. Additionally, in [14], a VDC-based separate meter-in separate meter-out controller has been presented to not only control the 3-DoF HHM but also improve the energy efficiency of the hydraulic actuator, resulting in a 47% reduction in energy consumption. In [15], an adaptive controller with gravity and friction identification has been developed for n-link hydraulic manipulators, and experimental results have been performed for 5-DoF HHM. Moreover, the combination of the time delay control with terminal SMC has been utilized to address the control problems of excavators in [16]. In both [17] and [18] the radial basic function neural networks (RBFNNs) are employed to estimate uncertainties.
@@ I-B3 Unmodeled uncertainties
Due to their universality and perfect capabilities in function approximation, RBFNNs are widely used in robotics and control applications to estimate model uncertainty and disturbances [29, 30]. In [17], a neural adaptive dynamic surface method has been proposed to control the n-link hydraulic manipulator, similar to the traditional backstepping method. The asymptotic convergence of errors with SGUUB has been proved for the system in the presence of model uncertainty and disturbance. A prescribed performance-based adaptive neural controller has been designed in [18] to achieve a good performance for HHM. In [31], uncertainties of hydraulic actuators are estimated using RBFNNs and UUB is proved with asymptotic tracking of the desired trajectory, highlighting the importance of achieving UUB in the presence of uncertainties. In the mentioned works, RBFNN is used to estimate disturbances and uncertainties in a rigid body-actuator model, while separately performing the estimations in actuator and rigid body parts will increase the accuracy. This is especially important in HHMs where the governing dynamics of the actuator and rigid body are completely different (the actuator is governed by fluid dynamics), and separate RBFNNs with corresponding input should be designed to learn the nature of the subsystem. It is stated in [32] that learning for complex systems (like uncertainty and disturbance estimation in 6-DoF HHM) should be solved using the decomposition and re-composition methods. Consequently, having a decentralized RBFNNs to deal with actuator and rigid body dynamics can boost the control performance and is of high importance.
@@ I-B4 VDC scheme
In addition to its better performance, the VDC scheme as a baseline controller offers other benefits, as follows: 1) It decomposes the entire complex system into subsystems where the decentralized controller and stability analysis can be performed; 2) the dynamics of each subsystem remains relatively simple, with fixed dynamic structures invariant to the target system; and 3) all uncertainties can be addressed at local subsystems. Moreover, the VDC controller designed for the electric manipulator can be applied to the hydraulic manipulator by only changing the subsystem related to the actuator while the rigid body part remains the same [12]. This is unlike other methods, in which applying a controller designed for electric robots to hydraulic robots requires redesigning the entire control scheme. Additionally, VDC has displayed a great performance in various fields, such as bilateral teleoperation [33], impedance control [34], and physical human-robot interaction [35]. All the mentioned benefits of VDC express the motivation of this paper to employ it as a baseline controller in this study.
@@ I-C1 Aim
The aim of the current paper is to achieve a robust and high precision control law for a generic 6-DoF HHMs consisting of anthropomorphic arm and spherical wrist. For doing so, the ORC scheme is proposed to consider compound input nonlinearities along with unmodeled uncertainties to achieve a robust and accurate tracking performance.
@@ I-C2 Contributions
Considering all the above-mentioned issues, compound input nonlinearities, unknown model uncertainties, and unknown disturbances are incoporated into control law to ensure the robustness and perfect performance of the controller. Altogether, this has given rise to the ORC scheme: orchestrating all the local controllers at the rigid body and actuator subsystems that are equipped to achieve high-precision control. As such, the contributions of this study can be stated with their significance as follows:
@@
  • Taking the presented ORC into account along with the complexity of the system, semi-global uniformly ultimate boundedness (SGUUB) is achieved for the first time for VDC. This is a significant contribution, since achieving such stability is mathematically challenging because of the existence of VPFs and the fact that VDC deals with each subsystem.
  • The results are experimentally validated using a 6-DoF commercial hydraulic manipulator, shown in Fig. 1. This is another significance of the presented paper. Although in [18, 17] neural adaptive controllers are designed for hydraulic manipulators, only simulation results of a 2-DoF robot were provided.
@@ II Mathematical preliminaries
In this section, the basic mathematical foundation of the VDC approach along with important lemmas and assumptions are presented, which are essential for control design and stability analysis. Fig. 1 demonstrates the schematics of the generic 6-DoF HHM with anothromorphic arm and spherical wrist. Fig. 1a shows the kinematic details with base angle ($\zeta_{1}$), lift angle ($\zeta_{2}$), tilt angle ($\zeta_{3}$), and spherical wrist angles of ($\xi_{1}$),($\xi_{2}$), and ($\xi_{3}$). Further, Fig. 1b displays the system that decomposed into three objects, each of which will be analyzed separately, by exploiting the virtual cutting point (VCP) concept of VDC.
@@
While deep neural networks (DNNs) can model complex patterns by stacking multiple layers, this often requires significant computational resources and careful tuning. Since inn real-time systems where low latency and fast processing are critical, RBFNNs are often a better choice than DNNs.
@@ V-B Hydraulic actuator dynamics and control
where $\mathcal{W}(u)$ is the selection function, defined as:
$\mathcal{W}(u)=\left\{\begin{array}[]{ll}1,&\text{if}\,\quad u>0,\\ 0,&\text{if}\,\quad u\leqslant 0,\\ \end{array}\right.$ (91)
and $\mathcal{Z}(\Delta p)$ is the function related to drops in pressure:
$\mathcal{Z}(\Delta p)=\text{sign}(\Delta p)\,\sqrt{\left\lvert\Delta p\right\rvert}.$ (92)
@@
where $Y_{v}=[Y_{v1},Y_{v2},Y_{v3},Y_{v4}]$ with $Y_{v1}=-\dfrac{\upsilon(p_{s}-p_{a})}{V_{0a}/A_{a}+x}\mathcal{S}(u)u$, $Y_{v2}=-\dfrac{\upsilon(p_{s}-p_{b})}{V_{0b}/A_{b}+(s-x))}\mathcal{S}(-u)u$, $Y_{v3}=-\dfrac{\upsilon(p_{a}-p_{r})}{V_{0a}/A_{a}+x}\mathcal{S}(-u)u$, and $Y_{v4}=-\dfrac{\upsilon(p_{b}-p_{r})}{V_{0b}/A_{b}+(s-x))}\mathcal{S}(u)u$, and $\theta_{v}=[c_{p1},c_{p2},c_{n1},c_{n2}]^{T}$.
@@
$\begin{split}u&=\dfrac{u_{f}}{c_{p1}\,\dfrac{\upsilon(p_{s}-p_{a})}{V_{0a}/A_{a}+x}+c_{n2}\dfrac{\upsilon(p_{b}-p_{r})}{V_{0b}/A_{b}+(s-x))}}\mathcal{S}(u_{f})\,\\ &+\dfrac{u_{f}}{c_{n1}\,\dfrac{\upsilon(p_{a}-p_{r})}{V_{0a}/A_{a}+x}+c_{p2}\dfrac{\upsilon(p_{s}-p_{b})}{V_{0b}/A_{b}+(s-x))}}\mathcal{S}(-u_{f}).\end{split}$ (98)
@@
$\begin{split}u&=\dfrac{u_{fr}\mathcal{S}(u_{fr})}{\hat{c}_{p1}\,\dfrac{\upsilon(p_{s}-p_{a})}{V_{0a}/A_{a}+x}+\hat{c}_{n2}\dfrac{\upsilon(p_{b}-p_{r})}{V_{0b}/A_{b}+(s-x))}}\,\\ &+\dfrac{u_{fr}\mathcal{S}(-u_{fr})}{\hat{c}_{n1}\,\dfrac{\upsilon(p_{a}-p_{r})}{V_{0a}/A_{a}+x}+\hat{c}_{p2}\dfrac{\upsilon(p_{s}-p_{b})}{V_{0b}/A_{b}+(s-x))}}.\end{split}$ (103)

### v4

# Orchestrated Robust Controller for Precision Control of Heavy-duty Hydraulic Manipulators

@@ Abstract
Vast industrial investment along with increased academic research on heavy-duty hydraulic manipulators has unavoidably paved the way for their automatization, necessitating the design of robust and high-precision controllers. In this study, an orchestrated robust controller is designed to address the mentioned issue for generic manipulators with an anthropomorphic arm and spherical wrist. Thanks to virtual decomposition control (VDC), the entire robotic system is decomposed into subsystems, and a robust controller is designed at each local subsystem by considering unknown model uncertainties, unknown disturbances, and compound input nonlinearities. As such, radial basic function neural networks (RBFNNs) are incorporated into VDC to tackle unknown disturbances and uncertainties, resulting in novel decentralized RBFNNs. All robust local controllers designed at each local subsystem, then, are orchestrated to accomplish high-precision control. In the end, for the first time in the context of VDC, a semi-globally uniformly ultimate boundedness is achieved under the designed controller. The validity of the theoretical results is verified by performing extensive simulations and experiments on a 6-degrees-of-freedom industrial manipulator with a nominal lifting capacity of $600\,kg$ at $5$ meters reach. Comparing the simulation result to the state-of-the-art controller along with provided experimental results, demonstrates that proposed method established all promises and performed excellently.
@@
{NtP}
Heavy-duty hydraulic manipulators (HHMs) play a pivotal role in various industrial applications, where their automation is critical for enhancing productivity, precision, and operational efficiency. However, automating HHMs presents significant challenges due to their complex dynamics, highly nonlinear behavior, and presence of unknown real-world uncertainties, such as unknown disturbances, non-ideal devices, unmodeled dynamics, and limited computational resources. These challenges underscore the need for control algorithms that are robust to such uncertainties, computationally efficient for real-time deployment, and capable of achieving high accuracy for reliable execution of free-motion tasks. This study designs a control scheme for generic 6-degree-of-freedom (DoF) HHMs, which ensures robustness and accuracy. Theoretical analysis along with experimental validations on a 6-DoF industrial HHM perfectly display stability of the method, robustness to real-world unknown uncertainties, efficiency in real-time implementation, and excellent accuracy in task execution. These results highlight scheme’s reliability and universality for industrial HHM applications, making it a significant contribution to advancing automation in this field. Additionally, detailed performance index provided with results enhances replicability and measurability, paving the way for meaningful comparisons in the future.
@@
SGUUB semi-global UUB
@@ I-A Heavy-duty Hydraulic Manipulators
Because of their higher robustness and remarkably larger payload-to-weight ratio (PWR) [1], heavy-duty hydraulic manipulators (HHMs) have been widely utilized in various fields, including construction, forestry, and agriculture. For example, the commercial manipulator shown in Fig. 1 has a lifting capacity of 600 kg at 5 meters reach, with a total weight of 445 kg (PWR almost 1:1). In contrast, the ABB robot [2] has a lifting capacity of 550 kg at 4.2 meters with the robot weight of 4,600 kg (PWR of 1:8). Such structural rigidity of the electrical manipulators is important to achieve higher accuracy in the manufacturing industry. However, in the field robotic applications, lightweight HHMs are particularly suited for installation on mobile machines, addressing the critical requirement for traversability. One prominent application of HHMs is in the mining industry, where their exceptional power is crucial for heavy-duty operations in challenging environments such as rock drilling [3]. Additionally, HHMs have offshore applications, where they are used for heavy-object manipulation in complex conditions. Consequently, it has become crucial to design a controller for HHMs that preserves robustness and precision in various industrial applications. The current vast amount of industrial investments with increased academic interests will revolutionize the HHMs industry [4, 5, 3, 6, 7, 8].
Despite their advantages, HHMs suffer from fundamental challenges, such as model nonlinearities and complexities emerging from structural complicatedness and fluid dynamics[8]. In contrast to electric actuators, whose model complexity can be neglected, hydraulic actuators are governed by fluid dynamics and may be subjected to nonsmooth and discontinuous nonlinearities due to friction between cylinder and piston, compound input nonlinearities, and valve under/overlap. Opposite to the electrically actuated open-chain manipulators, HHMs mostly comprise parallel-serial structures [9] that convert linear piston motion to rotational motion. Meanwhile, in rotary hydraulic actuators, the helical gears are employed to convert the piston motion to rotation, the same mechanism in the wrist actuators of the manipulator in this study (Fig. 1), each of which can deliver $2,000\,Nm$ of torque. The mentioned nonlinearities and complexities render the stability analysis and control design of HHMs significantly challenging, emphasizing the need for robust and high-performance control schemes, not only in theory but also in real-world applications.
@@ I-B1 Model-based controller
In [11], different NMBC methods, such as sliding mode control (SMC), adaptive inverse dynamics controller (AIDC), and model-reference adaptive controllers with velocity measurement (MRACV), have been designed and compared. The results unveil the better performance of MRACV in comparison to AIDC and SMC. In [12], an observer-based adaptive robust controller (ARC) has been designed to control a hydraulic manipulator, in which the backstepping approach has been employed to derive the control law. However, the experimental results in [13] with desired compensation ARC (DCARC) have shown that DCARC yields lower tracking error than ARC. Further, in [14] an adaptive output force tracking controller has been designed to control a 6-degrees-of-freedom (DoF) hydraulic manipulator by employing the virtual decomposition control (VDC) scheme, which was first introduced in [15]. A VDC control, also, in [16] has been designed for the precision control of a hydraulic crane. Additionally, in [17], a VDC-based separate meter-in separate meter-out controller has been presented to not only control the 3-DoF HHM but also improve the energy efficiency of the hydraulic actuator, resulting in a 47% reduction in energy consumption. In [18], an adaptive controller with gravity and friction identification has been developed for n-link hydraulic manipulators, and experimental results have been performed for 5-DoF HHM. Moreover, the combination of the time delay control with terminal SMC has been utilized to address the control problems of excavators in [19]. In both [20] and [21] the radial basic function neural networks (RBFNNs) are employed to estimate uncertainties. Additionally, in [22] and [23] finite time tracking control and model-data hybrid control are proposed to control hydraulic manipulators, respectively.
@@ I-B3 Unmodeled uncertainties
Due to their universality and perfect capabilities in function approximation, RBFNNs are widely used in robotics and control applications to estimate model uncertainty and disturbances [34, 35]. In [20], a neural adaptive dynamic surface method has been proposed to control the n-link hydraulic manipulator, similar to the traditional backstepping method with the asymptotic convergence of errors. A prescribed performance-based adaptive neural controller has been designed in [21] to achieve a good performance for HHM. In [36], uncertainties of hydraulic actuators are estimated using RBFNNs and UUB is proved with asymptotic tracking of the desired trajectory, highlighting the importance of achieving UUB in the presence of uncertainties. In the mentioned works, RBFNN is used to estimate disturbances and uncertainties in a rigid body-actuator model, while separately performing the estimations in actuator and rigid body parts will increase the accuracy. This is especially important in HHMs where the governing dynamics of the actuator and rigid body are completely different (the actuator is governed by fluid dynamics), and separate RBFNNs with corresponding input should be designed to learn the nature of the subsystem. It is stated in [37] that learning for complex systems (like uncertainty and disturbance estimation in 6-DoF HHM) should be solved using the decomposition and re-composition methods. Consequently, having a decentralized RBFNNs to deal with actuator and rigid body dynamics can boost the control performance and is of high importance.
@@ I-B4 VDC scheme
In addition to its better performance, the VDC scheme as a baseline controller offers other benefits, as follows: 1) It decomposes the entire complex system into subsystems where the decentralized controller and stability analysis can be performed; 2) the dynamics of each subsystem remains relatively simple, with fixed dynamic structures invariant to the target system; and 3) all uncertainties can be addressed at local subsystems. Moreover, the VDC controller designed for the electric manipulator can be applied to the hydraulic manipulator by only changing the subsystem related to the actuator while the rigid body part remains the same [15]. This is unlike other methods, in which applying a controller designed for electric robots to hydraulic robots requires redesigning the entire control scheme. Additionally, VDC has displayed a great performance in various fields, such as bilateral teleoperation [38], physical human-robot interaction [39], and impedance control [40]. All the mentioned benefits of VDC express the motivation of this paper to employ it as a baseline controller in this study.
@@ I-C1 Aim
The aim of the current paper is to achieve a robust and high precision control law for generic 6-DoF HHMs consisting of anthropomorphic arm and spherical wrist. For doing so, the ORC scheme is proposed to consider compound input nonlinearities along with unmodeled uncertainties to achieve a robust and accurate tracking performance.
@@ I-C2 Contributions
Considering all the above-mentioned issues, compound input nonlinearities, unknown model uncertainties, and unknown disturbances are incorporated into control law to ensure the robustness and perfect performance of the controller. Altogether, this has given rise to the ORC scheme: orchestrating all the local controllers at the rigid body and actuator subsystems that are equipped to achieve high-precision control. As such, the contributions of this study can be stated with their significance as follows:
@@
  • Taking the presented ORC into account along with the complexity of the system, semi-global uniformly ultimate boundedness (SGUUB) is achieved for the first time for VDC. Establishing such stability in the context of VDC is mathematically challenging; however, it is a crucial step in ensuring the robustness of the proposed method for real-world applications, marking another significant contribution.
  • The results are experimentally validated using a 6-DoF commercial hydraulic manipulator, shown in Fig. 1. This is another significance of the presented paper, which showcases the reliability and universality of the proposed method in real-world industrial application.
@@ II Mathematical preliminaries
In this section, the basic mathematical foundation of the VDC approach along with important lemmas and assumptions are presented, which are essential for control design and stability analysis. Fig. 1 demonstrates the schematics of the generic 6-DoF HHMs with anothromorphic arm and spherical wrist. Fig. 1a shows the kinematic details with base angle ($\zeta_{1}$), lift angle ($\zeta_{2}$), tilt angle ($\zeta_{3}$), and spherical wrist angles of ($\xi_{1}$),($\xi_{2}$), and ($\xi_{3}$). Further, Fig. 1b displays the system that decomposed into three objects, each of which will be analyzed separately, by exploiting the virtual cutting point (VCP) concept of VDC.
@@
While deep neural networks (DNNs) can model complex patterns by stacking multiple layers, this often requires significant computational resources and careful tuning. Since in real-time systems where low latency and fast processing are critical, RBFNNs are often a better choice than DNNs.
@@ III-B3 Dynamics of Object 1
Remark 2
The equations in (8)-(30) represent the kinematics of the entire system, which is decomposed into subsystems, as shown in Figs. 2-4. These equations are required for force analysis in (33)-(45), which are the representation of the system dynamics. The velocities and forces of each subsequent frame are computed through the transformation matrix defined in (2). This approach ensures that the analysis is grounded in precise, systematic computations, facilitating the accurate modeling of the manipulator’s behavior.
@@ IV-C Required Force Vectors
Building upon the system’s kinematics and dynamics analyses presented in Section III, this section proposes a robust control algorithm designed to achieve the system’s objectives effectively. The equations in (48)-(62) describe the expected kinematic behavior of the system when following the desired commands, laying the foundation for our control strategy. Leveraging this model-informed data, we then calculate the necessary forces to enforce the desired system behavior as in (66)-(74). Furthermore, the proposed algorithm accounts for unknown model uncertainties by incorporating RBFNNs as detailed in (65), ensuring robustness in the face of unpredictable real-world conditions.
Remark 4
@@ V-A Inverse Deadzone-Backlash
Remark 5
As demonstrated, the complex and highly nonlinear nature of the compound input nonlinearities, as described in (77), is transformed into a more manageable error dynamic representation in (83). This compact formulation enables the incorporation of these challenging nonlinearities directly into the control law, allowing for effective handling of them to ensure both robustness and precision in system performance.
@@ V-B Hydraulic actuator dynamics and control
where $\mathcal{W}(\mathrm{v})$ is the selection function for the given signal $\mathrm{v}$, defined as:
$\mathcal{W}(\mathrm{v})=\left\{\begin{array}[]{ll}1,&\text{if}\,\quad\mathrm{v}>0,\\ 0,&\text{if}\,\quad\mathrm{v}\leqslant 0,\\ \end{array}\right.$ (91)
and $\mathcal{Z}(\cdot)$ is the function related to drops in pressure:
$\mathcal{Z}(\cdot)=\text{sign}(\cdot)\,\sqrt{\left\lvert\cdot\right\rvert}.$ (92)
@@
where $Y_{v}=[Y_{v1},Y_{v2},Y_{v3},Y_{v4}]$ with $Y_{v1}=-\dfrac{\mathcal{Z}(p_{s}-p_{a})}{V_{0a}/A_{a}+x}\mathcal{W}(u)u$, $Y_{v2}=-\dfrac{\mathcal{Z}(p_{s}-p_{b})}{V_{0b}/A_{b}+(s-x))}\mathcal{W}(-u)u$, $Y_{v3}=-\dfrac{\mathcal{Z}(p_{a}-p_{r})}{V_{0a}/A_{a}+x}\mathcal{W}(-u)u$, and $Y_{v4}=-\dfrac{\mathcal{Z}(p_{b}-p_{r})}{V_{0b}/A_{b}+(s-x))}\mathcal{W}(u)u$, and $\theta_{v}=[c_{p1},c_{p2},c_{n1},c_{n2}]^{T}$.
@@
$\begin{split}u&=\dfrac{u_{f}}{c_{p1}\,\dfrac{\mathcal{Z}(p_{s}-p_{a})}{V_{0a}/A_{a}+x}+c_{n2}\dfrac{\mathcal{Z}(p_{b}-p_{r})}{V_{0b}/A_{b}+(s-x))}}\mathcal{W}(u_{f})\,\\ &+\dfrac{u_{f}}{c_{n1}\,\dfrac{\mathcal{Z}(p_{a}-p_{r})}{V_{0a}/A_{a}+x}+c_{p2}\dfrac{\mathcal{Z}(p_{s}-p_{b})}{V_{0b}/A_{b}+(s-x))}}\mathcal{W}(-u_{f}).\end{split}$ (98)
@@
$\begin{split}u&=\dfrac{u_{fr}\mathcal{W}(u_{fr})}{\hat{c}_{p1}\,\dfrac{\mathcal{Z}(p_{s}-p_{a})}{V_{0a}/A_{a}+x}+\hat{c}_{n2}\dfrac{\mathcal{Z}(p_{b}-p_{r})}{V_{0b}/A_{b}+(s-x))}}\,\\ &+\dfrac{u_{fr}\mathcal{W}(-u_{fr})}{\hat{c}_{n1}\,\dfrac{\mathcal{Z}(p_{a}-p_{r})}{V_{0a}/A_{a}+x}+\hat{c}_{p2}\dfrac{\mathcal{Z}(p_{s}-p_{b})}{V_{0b}/A_{b}+(s-x))}}.\end{split}$ (103)